OpenStudy (anonymous):

Simplify: (sin Θ − cos Θ)2 + (sin Θ + cos Θ)2 1 2 sin2 Θ cos2 Θ @Gin_Ichimaru @dan815 @zuberbier120 @amistre64 @love_jessika15 @YanaSidlinskiy

OpenStudy (anonymous):

2

OpenStudy (anonymous):

2

OpenStudy (solomonzelman):

(sin Θ − cos Θ)² + (sin Θ + cos Θ)² applying ` (a-b)²=a²-2ab+b² ` and ` (a+b)²=a²+2ab+b²` sin²Θ + 2sinΘcosΘ− cos Θ² + sin²Θ - 2sinΘcosΘ− cos Θ² 2sin²Θ − 2cos Θ²

OpenStudy (solomonzelman):

My bad, yes it is a 2

OpenStudy (solomonzelman):

the are positive. sin²Θ + 2sinΘcosΘ+ cos Θ² + sin²Θ - 2sinΘcosΘ+ cos Θ²

OpenStudy (anonymous):

ok thank you

OpenStudy (solomonzelman):

the only thing is that you just posted a plain answer, but other than that...

OpenStudy (anonymous):

its two don't let him run u in circles she just wants the answer

OpenStudy (solomonzelman):

Good work, but if someone wants just the answer, it is not a reason to violate the rule.

OpenStudy (solomonzelman):

\[\prod_{ }^{ }\sum_{ }^{ }\Huge{\times }t~~~ti me,~~~I'l l ~~~report.\]

OpenStudy (anonymous):

what rule u nerd

OpenStudy (yanasidlinskiy):

Calling someone a nerd? You've agreed to the CoC not to be rude:)

OpenStudy (anonymous):

coc??

OpenStudy (yanasidlinskiy):

http://openstudy.com/code-of-conduct